OpenStudy (anonymous):

The price of an extra-large pizza increased from $12 to $15. What was the percent of increase in the cost of the pizza? Round to the nearest percent, if necessary. A. 25% B. 3% C. 20% D. 15%

OpenStudy (johnweldon1993):

Started at 12...went to 15...that's a 3 dollar change So what percent of the original 12...is 3? \[\large \frac{3}{12} = \frac{1}{4} = ?\]

OpenStudy (anonymous):

it goes from twelve dollars to 15 so thats a $3/$12 increase thus making it 1/4 and 25%

OpenStudy (anonymous):

25%

OpenStudy (anonymous):

@chakradhar you need to show calculations not just plain answers.

OpenStudy (anonymous):

yes..increase in cost=15-12=3$..so,(3/initialcost)=3/12 gives u the increased cost of pizza over initial cost..and 2 get % ,just multiply by 100.